Hi could you please solve this problem for me. I don't understand how to set it up. Helped by a strong jet stream, a Los-Angeles-to-Boston plane flew 10% faster than usual and made the 4400km trip in 30min less time then usual. At what speed does the plane usually fly?
What i came up with:
x=old speed
x-30=new speed?

Answer by venugopalramana(3286)

Hi could you please solve this problem for me. I don't understand how to set it up. Helped by a strong jet stream, a Los-Angeles-to-Boston plane flew 10% faster than usual and made the 4400km trip in 30min less time then usual. At what speed does the plane usually fly?
What i came up with:
x=old speed...OK
x-30=new speed?..NO 30 MINUTES LESS IS TIME TAKEN.NOT SPEED.HE SAID SPEED IS 10% FASTER,SO IF OLD SPEED IS X
NEW SPEED IS FASTER BY =X*10/100=0.1X
SO NEW SPEED =X+0.1X=1.1X
SO NOW TIME TAKEN AT NEW SPEED =DISTANCE/SPEED=4400/(1.1X)
TIME TAKEN AT OLD SPEED =4400/X
DIFFERENCE = 30 MTS=30/60 HRS=0.5 HRS.
SO WE GET
4400/X - 4400//1.1X = 0.5
4400{1/X - 1/1.1X)=0.5
4400(1.1-1)/1.1X=0.5
4400(0.1/1.1X)=0.5
4400*1/11X=0.5
11X*0.5=4400
X=4400/5.5=800 KMPH